2. In 10 years, a $200,000,000 liability is due. Only one bond is available to you: ABC company, 10 years semiannual payments, 4% YTM, 8% coupon, 3% reinvestment rate, BBB credit rating.
a. Provide the # of bonds needed to fund the liability and the cost.
b. If the reinvestment rate falls from 3% to 2%, what is the projected deficit?
c. To prevent a deficit, explain how immunization can work, how it can help you and what conditions are required.
1
Expert's answer
2020-10-29T18:22:39-0400

Answer 1 a. Coupon Payment = Principal*rate of interest/2 = 1000*5.1%/2 = $ 25.5

b. Time line with each period of 6 months (total period = 20*2 = 40)

ANswer 2

From the diagram given in question

a. Maturity of bonds (in Years) = 50/2 25 years (divided by 2 because semi annual payments)

b. Coupion rate as a percentage = Total coupon interest in a year / Face vale * 100 = 19.92*2 / 1000 *100 = 3.984%

c. Face value = Maturity Value = $ 1,000